OpenStudy (anonymous):

PLEASE HURRY THERE'S A DEADLINE FOR IT IN 5 MINUTES About 8 to 10 U.S Households have TV remote controls. There were 129,200,000 U.S. Households in 1998. How many households have TV remote controls? SET UP A PROPORTION

OpenStudy (anonymous):

This is the proportion \[\frac{ 10 }{ 129,200,000 }=\frac{ 8 }{ ? }\] This is how to solve it. One of the most popular ways is by cross multiplying. So first 8 x 129,200,000 8 x 129,200,000 = 1,033,600,000 Next divide your answer by 10 1,033,600,000/10 = 103,360,000 So the answer is 103,360,000 households had remotes.

OpenStudy (anonymous):

SO IT WOULD BE 8/103,360,000 OR JUST 103,360,000

OpenStudy (anonymous):

depending how they want it written, it would either be 8/1,033,600,000 or just 103,360,000.